The half life of a specific Tc-99m is 6 hours. How much of an 192 mg sample will have decayed after 30 hours? 186 mg 32 mg 6 mg 160 mg The answer is : 186 mg. Why?
Answer #1

After every half life (6hrs) number of radioactive nuclide decrease to its half value.

Number of nuclide is proportional to mass of radio active material. Let initially there were No active nuclides.

Concentration = Co/2n where n = number of half lives

= time /t50%

After 6 hours radi active nuclide becomes No/2

After 12 hours radio active nuclide becomes No/22

After 18 hours radio active nuclide becomes No​​​​​​/23

After 24 hours radio active nuclide becomes No/24

After 30 hours radio active nuclide becomes No/25

No =( 192×10-3/(atomic mass))Na where Na is Avogadro number.

N = No/ 32

Mass of sample left after 30 hours = 192÷32 = 6mg

Mass of sample decayed after 30 hours = 192-6 = 186mg